OpenStudy (anonymous):

f a light flashes every 6second, how many times it flash in 3/4 of an hours?

OpenStudy (anonymous):

Firstly find what is 3/4 of an hour...

OpenStudy (anonymous):

\[\frac{3}{4} \times 60 = ?? (Minutes)\]

ganeshie8 (ganeshie8):

\(\large \frac{3}{4} hour \times \frac {60 \ minutes} {1\ hour} \times \frac{ 60\ seconds}{1\ minute} \times \frac{1 \ flash}{6 \ seconds }\)